Please use the Transfers Page for information on data transfer.
See the Modules Page
Please enter any significant events related to the AuScope VLBI project, including observations, configuration changes, hardware upgrades etc.
Three major repairs were finished on 2/12/15
* DBBC * IF wiring * Anemometer
DBBC: The Katherine DBBC is fully repaired as of 2/12/15. It originally failed 24/11/15, during r1715 with a total comms failure. This was eventually traced to the PCI7200 board, where the driver chip had burnt out. Replacing this board allowed Katherine to resume observations but with some apparent communication problems with the core2 units - bbc attenuations were being mis-set dur to a stuck bit in the communications line. The problem developed further after an initial inspection and no communications with either the CoMos or Core power levels could be made. The problem was eventually resolved on 1/12/15 when the PCI9111 board was replaced, but it may be that it was a connection issue instead - the PC frame is warped at Ke and it's possible to have some PCI lnes now connected through the backplane to the CPU.
IF wiring: The power level of SRCP going into CoMo D was unusually low in testing after the repair. It appears that the hardline or SMA connector was not making a good connection between the IF box and DBBC. Removing a replacing the hardline between the IF box output and DBBC radically improved the power level (~30 dB).
Anemometer: The anemometer at Katherine has suffered a fault and is reporting 1000 km/h winds. It has been turned off (using ispke) to avoid triggering spurious windstows. Until it is repaired, observers should use the weather radar at Katherine to look out for storm activity at Katherine.
The S/X was replaced on Hb12 on 13/11/15. Initial tests with Y-factor tests indicate good X-band Tsys (~78 K) but fairly poor S-band Tsys (95-105 K). Pointing data was collected over the weekend and a new pointing model was generated (2015.317) but the corrections are small.
A new version of the IFBox server has been written and installed at Hb, Ke and Yg. It should be more robust than the previous version and reduce the number of spurious “odd response from the pic” messages. The new software is /usr2/AuscopeUtils/monitoring/PIC_interfaces/Monica.IFBox.Sep2015.py - the /etc/init.d/Monica.IFBox scripts have been updated to use this from now on.
The list of NTP servers has been updated on the field system and Mk5 computers. All have had the following servers added:
0.pool.ntp.org 1.pool.ntp.org 2.pool.ntp.org 3.pool.ntp.org
su vi /etc/ntp.conf (added the listed servers) /etc/init.d/ntp restart
The log monitor has had a major update. New features:
If you experience any problems with the new version, the old one can be used by typing pcfs_log_monitor_old.pl at the ops2 prompt.